OpenStudy (anonymous):

4x - 3y = 10 make y the subject of the formula

OpenStudy (radar):

-3y=-4x+10 y=(4/3)x - 10/3

OpenStudy (radar):

I hope this what is being asked!

OpenStudy (anonymous):

i got a negative 3 though...

OpenStudy (anonymous):

i got -3y =10-4x

OpenStudy (anonymous):

then i finished off with y = (10-4x) / -3